# They See Me Rolling: High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Inspection

### v2

@@ III Background - Event Camera Principles
Unlike conventional cameras that capture intensity information across all pixels at fixed frame rates, event cameras employ a fundamentally distinct, bio-inspired sensing paradigm [Gallego2022]. Each pixel operates independently and asynchronously, responding dynamically to changes in illumination. Specifically, a pixel monitors the logarithm of the incident light intensity, $L(x,y,t)=\log(I(x,y,t))$. An event $e_{k}=(x_{k},y_{k},t_{k},p_{k})$ is generated at coordinates $(x_{k},y_{k})$ only when the change in this logarithmic intensity, $\Delta L$, relative to its value at the time of the last event from that pixel, surpasses a predefined contrast threshold, $\pm C$. The generated event includes a high-resolution timestamp $t_{k}$ and a polarity $p_{k}\in\{-1,+1\}$, signifying the direction of the brightness change (decrease or increase, respectively). Following event generation, the pixel’s internal reference brightness level is updated, priming it for subsequent intensity variations. The output is thus not a sequence of frames, but a sparse, continuous stream of asynchronous events reflecting dynamic scene changes. The unique characteristics of event cameras offer substantial benefits for high-speed tactile inspection. Their inherent high temporal resolution, with events timestamped often at microsecond precision, allows for tactile sensing at high speeds without the motion blur experienced by other roller tactile sensors at similar speeds. Furthermore, event-based cameras benefit from low latency and high dynamic range allowing for potential reduction in computational load and robustness to varying illumination within the sensor housing respectively.
@@ IV-A Mechanical Design and Fabrication
Refer to caption 1 Side Connector    2 Side Cover    3 Bearing    4 Rotating Shaft    5 Top Connector    6 Camera Holder    7 DVXplorer mini camera    8 LED Ring with Diffuser    9 Acrylic Cylinder and Elastomer
Figure 2: Mechanical Design of Roller: Exploded view of roller components with roller dimensions and insert of assembled view.
@@ IV-B1 Event-Based Multi-View Stereo (EMVS)
In motion parallax, objects closer to the observer (or camera) appear to move faster across the field of view than objects farther away. This effect is due to the relative difference in angular displacement caused by motion, and it’s a key depth cue used by both biological and artificial vision systems to perceive 3D structure from motion. To estimate the 3D structure of the scanned surface from the event stream generated by the moving roller, we adapt the multi-view stereo (MVS) space-sweep method, originally developed for frame-based cameras [collins_1996_a] and later extended to event cameras (EMVS) [rebecq_2017_emvs]. This technique leverages the known viewpoints (poses) of the event camera over time. The core idea involves discretizing a volume of interest, $\bar{D}\in\mathbb{R}^{w\times h\times N_{z}}$, encompassing the potential 3D scene, where $w$, $h$, and $N_{z}$ represent the dimensions of the volume. Each captured event, $e_{k}=(x_{k},y_{k},t_{k},p_{k})$, is back-projected through this volume as a 3D ray using the camera’s calibrated intrinsic parameters and its known pose at time $t_{k}$. For our purposes, the physical design of the roller sensor provides a strong prior on the depth range at which events occur. We exploit this to optimize the discretization of the volume $\bar{D}$ for both accuracy and computational efficiency. The parameters $w$ and $h$ are set to the camera’s resolution of 640x480 respectively. Furthermore, the depth dimension is constrained to a 4 mm range, from a minimum depth of 43 mm to a maximum depth of 47 mm, corresponding to the expected range of elastomer deformation from the sensor’s 47 mm radius. This volume is discretized into $N_{z}=$ 500 depth planes, yielding a high depth resolution of 8 µm per plane. This constrained search space is critical for achieving both high-speed processing and high-precision reconstruction.
@@
A Disparity Space Image (DSI) is employed to accumulate evidence for surfaces within the volume. The DSI acts as a 3D histogram where each voxel counts the number of back-projected event rays passing through it, typically referenced to a specific viewpoint. Voxels corresponding to real surfaces in the scene are expected to lie at the intersection of rays from multiple viewpoints, resulting in high ray counts. Local maxima in the ray density within the DSI are then identified as potential 3D points belonging to the scanned surface [rebecq_2017_emvs, ayyad_2023_neuromorphic]. Processing these maxima using Adaptive Gaussian Thresholding, yields a semi-dense depth map relative to the reference viewpoint.
@@
Our implementation, in line with the original EMVS framework [rebecq_2017_emvs], conceptually associates a unique pose with every event. In our hardware setup (refer to Section V), the sensor pose is derived from the UR10 robot’s end-effector state, which is logged at 100 Hz. To obtain a pose for any event’s specific microsecond-resolution timestamp, we perform linear interpolation between the two nearest pose readings. For computational efficiency, as is done in EMVS, events are grouped into small temporal packets (comprised of 256 events) and all events in a packet are assigned the interpolated pose of the median event’s timestamp.
@@
where $z$ is the depth $Z_{C}$ of the point in the camera frame, $\mathbf{K}$ is the $3\times 3$ camera intrinsic matrix, $\mathbf{R}|\mathbf{t}]$ represents the $3\times 4$ extrinsic matrix (rotation $\mathbf{R}$, which is a $3\times 3$ matrix and translation $\mathbf{t}$, which is $3\times 1$) transforming world coordinates to camera coordinates $\mathbf{P}_{C}=[X_{C},Y_{C},Z_{C}]^{T}$. The absolute pose of the camera frame relative to the robot base frame, $^{B}\mathbf{T}_{C}(t)$, which defines $\mathbf{R}$ and $\mathbf{t}$ relative to the base frame at time $t$, is obtained by combining the robot’s end-effector pose $^{B}\mathbf{T}_{E}(t)$ (derived from forward kinematics) with the fixed hand-eye transformation $^{E}\mathbf{T}_{C}$, determined via hand-eye geometric calibration [ayyad_2023_neuromorphic]. The intrinsic matrix $\mathbf{K}$ is also obtained through calibration using E-Calib, a calibration toolbox for event cameras [salah_2024_ecalib].
@@ IV-B2 Multi-Reference Depth Map Generation and Fusion
However, roller-based tactile sensing can suffer from errors due to the sensor’s curved nature, leading to inconsistent contact depth and potential fabrication variations [li_2025_cyclic]. The curved elastomer means that even when contacting a flat surface, the contact points have varying depths relative to the camera. Second, the small contact area means any given surface feature is only observed over a short rolling distance within a short time window, resulting in a small baseline for multi-view stereo, which can make depth estimation more challenging. High-speed motion can also introduce occlusion and warping artifacts. To enhance robustness against these effects and specifically mitigate the depth inconsistency and curvature-induced errors inherent in cylindrical tactile sensing, we extend the EMVS framework by generating and fusing depth maps from multiple reference times within the same time window.
Inspired by works demonstrating the benefits of multi-reference objectives in event-based vision [shiba_2024_secrets] and fusion techniques in roller tactile sensing [li_2025_cyclic], our approach extends the EMVS framework through a three-stage fusion process. First, we generate three separate DSIs using the same 20 ms stream of events and their associated poses. Each DSI is constructed relative to a different reference pose: one at the start ($t_{s}$), one at the midpoint ($t_{m}$), and one at the end ($t_{e}$) of the time window. Although the input data is identical, the differing reference frames result in distinct DSI geometries and consequently, slightly different ray count distributions. Second, as is done in [rebecq_2017_emvs], for each of the three DSIs, we extract a deterministic depth map by performing a maximization along each pixel’s viewing ray to find the depth plane with the highest vote count. This yields three independent depth maps: $Z_{s}$, $Z_{m}$, and $Z_{e}$ referenced at the start ($t_{s}$), midpoint ($t_{m}$), and end ($t_{e}$) times of the window, respectively.
Third, we fuse these individual depth maps using a Bayesian Model Averaging (BMA) strategy (Figure 3), which has proven effective for combining multiple noisy estimates in vision tasks [sun_2018_bayesian]. This process averages out noise and inconsistencies present in the individual reconstructions, mitigating both DSI discretization artifacts and the impact of noise and occlusions. The final fused depth map, $Z_{f}$, is computed as a weighted average of the three deterministic maps after warping them to a common reference frame (chosen as the midpoint view, $t_{m}$):
@@ IV-C Parameter Calibration
To optimize the final reconstruction quality, we perform a one-time calibration using an object with precisely known geometry, specifically a sphere of known radius $R_{sphere}$. This calibration determines the optimal values for both the adaptive thresholding parameters used in processing the DSI and the fusion weights used in combining multi-reference depth maps. The adaptive thresholding parameters (constant $C$ and median filter size $G_{f}$) control the density and noise level of the individual depth maps $Z_{i}$, while the fusion weights $w_{s},w_{m},w_{e}$ dictate how these maps are combined into the final estimate $Z_{f}$. Furthermore, this calibration procedure can be readily repeated in-situ, enabling the system to adapt to gradual changes in the elastomer surface, such as wear, by re-optimizing these parameters.
To establish the ground truth depth map $Z_{gt}$ for the sphere contact, we adapt the methodology used in [zhang_gelroller_2024]. For calibration, the sensor is rolled over the calibration sphere while recording event data. The sphere is mounted in a fixed position on an optical breadboard such that the sphere passes directly under the camera’s field of view during rolling, ensuring consistent alignment (refer to setup in Figure 6). Then, the resulting circular contact patch in the depth map is identified (visualized in Figure 4). From this, we determine the radius of the pressed circle, $r$ (in pixels), and the coordinates of its center, $(x_{b},y_{b})$ (in pixels). The physical radius of the calibration sphere, $R_{sphere}$, is known in millimeters. The measured contact radius, $r$, (in pixels) is then converted to millimeters by multiplying by the scale factor $f$ (mm/pixel), where $f=0.0422$ mm/pixel. This scale factor was determined by imaging the gap between the jaws of a Vernier caliper indented directly on the sensor surface at known separations. Two readings were taken at each of four separations (2 mm, 3 mm, 4 mm, and 5 mm), the scale factor was computed for each reading, and the eight resulting values were averaged to obtain the final scale factor $f$. The depth of the sphere’s center, $z_{b}$ (in mm), can then be calculated geometrically as:
$z_{b}=-\sqrt{R_{sphere}^{2}-(r\cdot f)^{2}}$ (6)
@@
$\big((x_{i}-x_{b})\cdot f\big)^{2}+\big((y_{i}-y_{b})\cdot f\big)^{2}+(z_{i}-z_{b})^{2}=R_{sphere}^{2}$ (7)
@@
With the ground truth $Z_{gt}$ established, both sets of parameters (thresholding and fusion weights) are optimized simultaneously by minimizing the Mean Absolute Error (MAE) between the final fused depth map $Z_{f}$ and $Z_{gt}$. A region mask corresponding to the sphere’s contact area is applied before calculating the MAE. The optimization is performed via a grid search over the combined parameter space:
@@
The calibrated values for all parameters are subsequently fixed and utilized for processing the data from all other scanned objects presented in this work. The effect of calibration on the processed depth map and point cloud are shown in Figure 5. From a Bayesian perspective, the optimal weights should ideally be inversely proportional to the uncertainty associated with each individual depth map $Z_{i}$, effectively giving more influence to more confident estimates.
@@ IV-D Braille Feature Recognition
To demonstrate the sensor’s capability for fine feature detection at high speed, a task complementary to 3D geometric reconstruction, we implemented a Braille character recognition pipeline. This task serves as a well-established benchmark in the tactile sensing community for evaluating high-speed spatial resolution, allowing for a direct comparison with state-of-the-art systems like that of Potdar et al. [potdar_high-speed_2024]. This method specifically evaluates the sensor’s ability to capture fine textural details directly from the event stream, a key aspect of tactile inspection that is distinct from geometric reconstruction. The process begins by applying motion compensation [Salah2023, Gallego_CMAX_2018] to the recorded event stream as the sensor rolls over the Braille plate. Given the constant linear velocity of the roller, $v_{roller}$, which is commanded to and precisely controlled by the UR10 robotic arm as described in the Experimental Setup (Sec. V), we can determine the corresponding optical flow $\mathbf{\dot{U}}=(\dot{u},\dot{v})$ of features moving across the image plane with $[u,v]$ as the pixel coordinates. The roller’s motion involves predominantly horizontal motion, therefore $\dot{v}\approx 0$ and $\dot{u}=\frac{-F}{Z}v_{roller}$, where $F$ is the focal length of the camera and $Z$ is the depth of the scene (assumed to be constant depth of 47 mm which is the radius of the roller with maximum deformation of 4.5 mm).
Without motion compensation, if we were to simply create a frame-like image by plotting all events recorded within a specific time window at their original pixel locations (a process called direct accumulation), the relative movement between the sensor and the Braille plate would cause significant motion blur. This occurs because events generated by a single point on the moving Braille surface would be spread across multiple pixels in the resulting image (as seen in Figure 12a). Using a long accumulation time integrates more events but exacerbates this blur as features traverse many pixels, whereas a very short accumulation time minimizes blur but often captures too few events to form a clear representation, leading to degraded features [Salah2023, Gallego_CMAX_2018]. To counteract this, we warp events with batches consisting of a constant number of events ($N_{e}$ = 20,000 events) at the initial timestamp of the event stream as the reference time, $t_{ref}$. For each event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ where $\mathbf{x}_{k}=(x_{k},y_{k})$, its warped position $\mathbf{x}^{\prime}_{k}$ is calculated as:
$\mathbf{x}^{\prime}_{k}=\mathbf{x}_{k}-\mathbf{\dot{U}}(t_{k}-t_{ref})$ (9)
A motion-compensated frame, or Image of Warped Events (IWE) $I_{warp}$, is generated by accumulating the counts of warped events into a 2D histogram (ignoring the polarities of each event):
$I_{warp}=\sum^{N_{e}}_{k}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k})$ (10)

### v3

@@ III Background - Event Camera Principles
Unlike conventional cameras that capture intensity information across all pixels at fixed frame rates, event cameras employ a fundamentally distinct, bio-inspired sensing paradigm [19]. Each pixel operates independently and asynchronously, responding dynamically to changes in illumination. Specifically, a pixel monitors the logarithm of the incident light intensity, $L(x,y,t)=\log(I(x,y,t))$. An event $e_{k}=(x_{k},y_{k},t_{k},p_{k})$ is generated at coordinates $(x_{k},y_{k})$ only when the change in this logarithmic intensity, $\Delta L$, relative to its value at the time of the last event from that pixel, surpasses a predefined contrast threshold, $\pm C_{\text{th}}$. The generated event includes a high-resolution timestamp $t_{k}$ and a polarity $p_{k}\in\{-1,+1\}$, signifying the direction of the brightness change (decrease or increase, respectively). Following event generation, the pixel’s internal reference brightness level is updated, priming it for subsequent intensity variations. The output is thus not a sequence of frames, but a sparse, continuous stream of asynchronous events reflecting dynamic scene changes. The unique characteristics of event cameras offer substantial benefits for high-speed tactile inspection. Their inherent high temporal resolution, with events timestamped often at microsecond precision, allows for tactile sensing at high speeds without the motion blur experienced by other roller tactile sensors at similar speeds. Furthermore, event-based cameras benefit from low latency and high dynamic range allowing for potential reduction in computational load and robustness to varying illumination within the sensor housing respectively.
@@ IV-A Mechanical Design and Fabrication
1
Side Connector    2 Side Cover    3 Bearing    4 Rotating Shaft    5 Top Connector    6 Camera Holder    7 DVXplorer mini camera    8 LED Ring with Diffuser    9 Acrylic Cylinder and Elastomer
Figure 2: Mechanical Design of Roller: Exploded view of roller components with roller dimensions and inset of assembled view.
@@ IV-B1 Event-Based Multi-View Stereo (EMVS)
In motion parallax, objects closer to the observer (or camera) appear to move faster across the field of view than objects farther away. This effect is due to the relative difference in angular displacement caused by motion, and it’s a key depth cue used by both biological and artificial vision systems to perceive 3D structure from motion. To estimate the 3D structure of the scanned surface from the event stream generated by the moving roller, we adapt the multi-view stereo (MVS) space-sweep method, originally developed for frame-based cameras [11] and later extended to event cameras (EMVS) [40]. This technique leverages the known viewpoints (poses) of the event camera over time. The core idea involves discretizing a volume of interest, $\bar{D}\in\mathbb{R}^{W\times H\times N_{z}}$, encompassing the potential 3D scene, where $W$, $H$, and $N_{z}$ represent the dimensions of the volume. Each captured event, $e_{k}=(x_{k},y_{k},t_{k},p_{k})$, is back-projected through this volume as a 3D ray using the camera’s calibrated intrinsic parameters and its known pose at time $t_{k}$. For our purposes, the physical design of the roller sensor provides a strong prior on the depth range at which events occur. We exploit this to optimize the discretization of the volume $\bar{D}$ for both accuracy and computational efficiency. The parameters $W$ and $H$ are set to the camera’s resolution of 640x480 respectively. Furthermore, the depth dimension is constrained to a 4 mm range, from a minimum depth of 43 mm to a maximum depth of 47 mm, corresponding to the expected range of elastomer deformation from the sensor’s 47 mm radius. This volume is discretized into $N_{z}=$ 500 depth planes, yielding a high depth resolution of 8 µm per plane. This constrained search space is critical for achieving both high-speed processing and high-precision reconstruction.
@@
A Disparity Space Image (DSI) is employed to accumulate evidence for surfaces within the volume. The DSI acts as a 3D histogram where each voxel counts the number of back-projected event rays passing through it, typically referenced to a specific viewpoint. Voxels corresponding to real surfaces in the scene are expected to lie at the intersection of rays from multiple viewpoints, resulting in high ray counts. Local maxima in the ray density within the DSI are then identified as potential 3D points belonging to the scanned surface [40, 5]. Processing these maxima using Adaptive Gaussian Thresholding, yields a semi-dense depth map relative to the reference viewpoint.
@@
Our implementation, in line with the original EMVS framework [40], conceptually associates a unique pose with every event. In our hardware setup (refer to Section V), the sensor pose is derived from the UR10 robot’s end-effector state, which is logged at 100 Hz. To obtain a pose for any event’s specific microsecond-resolution timestamp, we perform linear interpolation between the two nearest pose readings. For computational efficiency, as is done in EMVS, events are grouped into small temporal packets (comprised of 256 events) and all events in a packet are assigned the interpolated pose of the median event’s timestamp.
@@
where $z$ is the depth $Z_{C}$ of the point in the camera frame, $\mathbf{K}$ is the $3\times 3$ camera intrinsic matrix, $\mathbf{R}|\mathbf{t}]$ represents the $3\times 4$ extrinsic matrix (rotation $\mathbf{R}$, which is a $3\times 3$ matrix and translation $\mathbf{t}$, which is $3\times 1$) transforming world coordinates to camera coordinates $\mathbf{P}_{C}=[X_{C},Y_{C},Z_{C}]^{T}$. The absolute pose of the camera frame relative to the robot base frame, $^{B}\mathbf{T}_{C}(t)$, which defines $\mathbf{R}$ and $\mathbf{t}$ relative to the base frame at time $t$, is obtained by combining the robot’s end-effector pose $^{B}\mathbf{T}_{E}(t)$ (derived from forward kinematics) with the fixed hand-eye transformation $^{E}\mathbf{T}_{C}$, determined via hand-eye geometric calibration [5]. The intrinsic matrix $\mathbf{K}$ is also obtained through calibration using E-Calib, a calibration toolbox for event cameras [43].
@@ IV-B2 Multi-Reference Depth Map Generation and Fusion
However, roller-based tactile sensing can suffer from errors due to the sensor’s curved nature, leading to inconsistent contact depth and potential fabrication variations [31]. The curved elastomer means that even when contacting a flat surface, the contact points have varying depths relative to the camera. Second, the small contact area means any given surface feature is only observed over a short rolling distance within a short time window, resulting in a small baseline for multi-view stereo, which can make depth estimation more challenging. High-speed motion can also introduce occlusion and warping artifacts. To enhance robustness against these effects and specifically mitigate the depth inconsistency and curvature-induced errors inherent in cylindrical tactile sensing, we extend the EMVS framework by generating and fusing depth maps from multiple reference times within the same time window.
Inspired by works demonstrating the benefits of multi-reference objectives in event-based vision [49] and fusion techniques in roller tactile sensing [31], our approach extends the EMVS framework through a three-stage fusion process. First, we generate three separate DSIs using the same 20 ms stream of events and their associated poses. Each DSI is constructed relative to a different reference pose: one at the start ($t_{s}$), one at the midpoint ($t_{m}$), and one at the end ($t_{e}$) of the time window. Although the input data is identical, the differing reference frames result in distinct DSI geometries and consequently, slightly different ray count distributions. Second, as is done in [40], for each of the three DSIs, we extract a deterministic depth map by performing a maximization along each pixel’s viewing ray to find the depth plane with the highest vote count. This yields three independent depth maps: $Z_{s}$, $Z_{m}$, and $Z_{e}$ referenced at the start ($t_{s}$), midpoint ($t_{m}$), and end ($t_{e}$) times of the window, respectively.
Third, we fuse these individual depth maps using a Bayesian Model Averaging (BMA) strategy (Figure 3), which has proven effective for combining multiple noisy estimates in vision tasks [51]. This process averages out noise and inconsistencies present in the individual reconstructions, mitigating both DSI discretization artifacts and the impact of noise and occlusions. The final fused depth map, $Z_{f}$, is computed as a weighted average of the three deterministic maps after warping them to a common reference frame (chosen as the midpoint view, $t_{m}$):
@@ IV-C Parameter Calibration
To optimize the final reconstruction quality, we perform a one-time calibration using an object with precisely known geometry, specifically a sphere of known radius $R_{\text{sphere}}$. This calibration determines the optimal values for both the adaptive thresholding parameters used in processing the DSI and the fusion weights used in combining multi-reference depth maps. The adaptive thresholding parameters (constant $C$ and median filter size $G_{f}$) control the density and noise level of the individual depth maps $Z_{i}$, while the fusion weights $w_{s},w_{m},w_{e}$ dictate how these maps are combined into the final estimate $Z_{f}$. Furthermore, this calibration procedure can be readily repeated in-situ, enabling the system to adapt to gradual changes in the elastomer surface, such as wear, by re-optimizing these parameters.
To establish the ground truth depth map $Z_{\text{gt}}$ for the sphere contact, we adapt the methodology used in [62]. For calibration, the sensor is rolled over the calibration sphere while recording event data. The sphere is mounted in a fixed position on an optical breadboard such that the sphere passes directly under the camera’s field of view during rolling, ensuring consistent alignment (refer to setup in Figure 6). Then, the resulting circular contact patch in the depth map is identified (visualized in Figure 4). From this, we determine the radius of the pressed circle, $r$ (in pixels), and the coordinates of its center, $(x_{b},y_{b})$ (in pixels). The physical radius of the calibration sphere, $R_{\text{sphere}}$, is known in millimeters. The measured contact radius, $r$, (in pixels) is then converted to millimeters by multiplying by the scale factor $f$ (mm/pixel), where $f=0.0422$ mm/pixel. This scale factor was determined by imaging the gap between the jaws of a Vernier caliper indented directly on the sensor surface at known separations. Two readings were taken at each of four separations (2 mm, 3 mm, 4 mm, and 5 mm), the scale factor was computed for each reading, and the eight resulting values were averaged to obtain the final scale factor $f$. The depth of the sphere’s center, $z_{b}$ (in mm), can then be calculated geometrically as:
$z_{b}=-\sqrt{R_{\text{sphere}}^{2}-(r\cdot f)^{2}}$ (6)
@@
$\big((x_{i}-x_{b})\cdot f\big)^{2}+\big((y_{i}-y_{b})\cdot f\big)^{2}+(z_{i}-z_{b})^{2}=R_{\text{sphere}}^{2}$ (7)
@@
With the ground truth $Z_{\text{gt}}$ established, both sets of parameters (thresholding and fusion weights) are optimized simultaneously by minimizing the Mean Absolute Error (MAE) between the final fused depth map $Z_{f}$ and $Z_{\text{gt}}$. A region mask corresponding to the sphere’s contact area is applied before calculating the MAE. The optimization is performed via a grid search over the combined parameter space:
@@
The calibrated values for all parameters are subsequently fixed and utilized for processing the data from all other scanned objects presented in this work. The effect of calibration on the processed depth map and point cloud is shown in Figure 5. From a Bayesian perspective, the optimal weights should ideally be inversely proportional to the uncertainty associated with each individual depth map $Z_{i}$, effectively giving more influence to more confident estimates.
@@ IV-D Braille Feature Recognition
To demonstrate the sensor’s capability for fine feature detection at high speed, a task complementary to 3D geometric reconstruction, we implemented a Braille character recognition pipeline. This task serves as a well-established benchmark in the tactile sensing community for evaluating high-speed spatial resolution, allowing for a direct comparison with state-of-the-art systems like that of Potdar et al. [38]. This method specifically evaluates the sensor’s ability to capture fine textural details directly from the event stream, a key aspect of tactile inspection that is distinct from geometric reconstruction. The process begins by applying motion compensation [44, 20] to the recorded event stream as the sensor rolls over the Braille plate. Given the constant linear velocity of the roller, $v_{\text{roller}}$, which is commanded to and precisely controlled by the UR10 robotic arm as described in the Experimental Setup (Sec. V), we can determine the corresponding optical flow $\mathbf{\dot{U}}=(\dot{u},\dot{v})$ of features moving across the image plane with $[u,v]$ as the pixel coordinates. The roller’s motion involves predominantly horizontal motion, therefore $\dot{v}\approx 0$ and $\dot{u}=\frac{-F}{Z}v_{\text{roller}}$, where $F$ is the focal length of the camera and $Z$ is the depth of the scene (assumed to be constant depth of 47 mm which is the radius of the roller with maximum deformation of 4.5 mm).
Without motion compensation, if we were to simply create a frame-like image by plotting all events recorded within a specific time window at their original pixel locations (a process called direct accumulation), the relative movement between the sensor and the Braille plate would cause significant motion blur. This occurs because events generated by a single point on the moving Braille surface would be spread across multiple pixels in the resulting image (as seen in Figure 12a). Using a long accumulation time integrates more events but exacerbates this blur as features traverse many pixels, whereas a very short accumulation time minimizes blur but often captures too few events to form a clear representation, leading to degraded features [44, 20]. To counteract this, we warp events with batches consisting of a constant number of events ($N_{e}$ = 20,000 events) at the initial timestamp of the event stream as the reference time, $t_{\text{ref}}$. For each event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ where $\mathbf{x}_{k}=(x_{k},y_{k})$, its warped position $\mathbf{x}^{\prime}_{k}$ is calculated as:
$\mathbf{x}^{\prime}_{k}=\mathbf{x}_{k}-\mathbf{\dot{U}}(t_{k}-t_{\text{ref}})$ (9)
A motion-compensated frame, or Image of Warped Events (IWE) $I_{\text{warp}}$, is generated by accumulating the counts of warped events into a 2D histogram (ignoring the polarities of each event):
$I_{\text{warp}}=\sum^{N_{e}}_{k}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k})$ (10)